OpenStudy (anonymous):

Simplify √98x^5 as a radical. if you could show me how to do this it'd be greatly appreciated

OpenStudy (anonymous):

sqrt(2*49*x*x^4) 7x^2*sqrt(2x)

OpenStudy (anonymous):

\[7x^2\sqrt{2x}\]

OpenStudy (anonymous):

\[\rightarrow\sqrt{98x^5} \implies \sqrt{49 \times 2 \times x^2 \times x^2 \times x} \implies \sqrt{\cancel{49} \times 2 \times \cancel{x^2} \times \cancel{x^2} \times x} \implies\]\[7\times x \times x \sqrt{2x} \implies 7x^2\sqrt{2x}\]

OpenStudy (anonymous):

@flowmaster1776 remember laws of exponents bro/sis..

OpenStudy (anonymous):

ok. thank you both! really helping me understand.

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

lazy to type like you.. @Calcmathlete lols..

OpenStudy (anonymous):

lol. \(\huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}\)

OpenStudy (anonymous):

lol how did you do that ?

OpenStudy (anonymous):

Here is the code I used without the \[ \] part. \huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}

OpenStudy (anonymous):

*without the \ [ \ ] part.

OpenStudy (anonymous):

huge LaTeX space color{red}{for} space color{blue}{the} space color{salmon}{win} color{orange}{!}

OpenStudy (anonymous):

\huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you'd have to put backward slahs followed by bracket on both ends.

OpenStudy (anonymous):

\[huge \]LaTeX \[space \]color{red}{for} \[space \]color{blue}{the} [\space \]color{salmon}{win} \[color{orange}{!}]

OpenStudy (anonymous):

\[\huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}\]

OpenStudy (anonymous):

lol i am idiot

OpenStudy (anonymous):

\[huge] \[LaTeX] \[space] \[color{red}{for}] \[space] \[color{blue}{the}] \[space] \[color{salmon}{win}] \[color{orange}{!}]

OpenStudy (anonymous):

:((((((

OpenStudy (anonymous):

heeelpppppp

OpenStudy (anonymous):

before \huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!} you have to put \[ in the front and \] at the end

OpenStudy (anonymous):

thats weird the symbols didn't show up...

OpenStudy (anonymous):

yeah. If you put them together as in \ [ and \ ] together, then it is the what initializes the \(\LaTeX\).

OpenStudy (anonymous):

[\huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}]

OpenStudy (anonymous):

Damn it !!

OpenStudy (anonymous):

Missing something...the \ at the end before the ]

OpenStudy (anonymous):

[\huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}\]

OpenStudy (anonymous):

what the hell

OpenStudy (anonymous):

you always put the slash before the bracket never bracket before slash

OpenStudy (anonymous):

you put the bracket before the slash at the beginning

OpenStudy (anonymous):

Yeah. \ [ \ ] in that order

OpenStudy (anonymous):

\[huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}\]

OpenStudy (anonymous):

Forgot the \ infront of the huge...

OpenStudy (anonymous):

\[\huge \LaTeX \space \color{red}{for} \space \color{blue}{the} \space \color{salmon}{win} \color{orange}{!}\]

OpenStudy (anonymous):

there ya go

OpenStudy (anonymous):

Atlast did it !! *sigh*

OpenStudy (anonymous):

Here...just for you :) \(Hidden \space\LaTeX \space group!\) http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)

OpenStudy (anonymous):

nice !! thanx

OpenStudy (anonymous):

np :)